The half-life of aspirin in your bloodstream is 12 hours. How long will it take for the aspirin to decay to 75% of the original dosage.
Answer by stanbon(75887)

The half-life of aspirin in your bloodstream is 12 hours. How long will it take for the aspirin to decay to 75% of the original dosage.
-------
0.75A = A(1/2)^(t/12)
-----------------
0.75 = (1/2)^(t/12)
-----------------------
t/12 = log(0.75)/log(1/2)
-----
t/12 = 0.4150
------------------------------------------
t = 4.9 hrs = 4 hr 54 min
----------------------------------------
Cheers,
Stan H.
----------------------